A market is described by the following supply-and-demand curves:
Qs=4p
Qd=400-p

The equilibrium price isand the equilibrium quantity is.
Suppose the government imposes a price ceiling of $60. This price ceiling is , and the market price will be. The quantity supplied will be, and the quantity demanded will be. Therefore, a price ceiling of $60 will result in .
Suppose the government imposes a price floor of $60. This price floor is , and the market price will be. The quantity supplied will beand the quantity demanded will be. Therefore, a price floor of $60 will result in .
Instead of a price control, the government levies a tax on producers of $10. As a result, the new supply curve is:
Qs=4(p-10)
1
Expert's answer
2017-07-31T12:10:07-0400
Qs = Qd; 4p=400-p; 5p = 400; p=$80; Qe=4*80 or 400-80 = 320
The equilibrium price is $80 and the equilibrium quantity is 320.
Suppose the government imposes a price ceiling of $60. This price ceiling is $60, and the market price will be $60. The quantity supplied will be Qs = 4*60 = 240, and the quantity demanded will be Qd=400-60=340. Therefore, a price ceiling of $60 will result in shortage.
Suppose the government imposes a price floor of $60. This price floor is $60, and the market price will be $80 (equilibrium price). The quantity supplied will be 4*80 = 320 and the quantity demanded will be 400-80=320. Therefore, a price floor of $60 don’t impact on the market price.
Instead of a price control, the government levies a tax on producers of $10. As a result, the new supply curve is:
Qs=4(p-10)
Qs=Qd; 4(p-10) = 400 – p; 4p-40 = 400-p; 5p=440; p=$88; Qe=4(88-10) or 400-88 = 312.
The equilibrium price is $88 and the equilibrium quantity is 312.